OpenStudy (anonymous):

Can someone please help me with Pre-Cal and show me the steps? The problem is: 14^(a+8.5)+3=9. I'm not sure if I should subtract 3 on both sides and if I do what steps do I take to figure out what a is equal to?

OpenStudy (anonymous):

The equation is \[14^{a+8.5}+3=9\]

OpenStudy (amistre64):

subtract, then log ...

OpenStudy (amistre64):

something, plus 3, is 9 ... we know that 6+3 = 9 so whatever a is, it has to make the term equal to 6 ...

OpenStudy (amistre64):

14^(a+k) = 6 log both sides ... log 14^(a+k) = log 6 now use log properties .... log p^n = n log p (a+k) log 14 = log 6 divide and subtract ...

OpenStudy (anonymous):

so even though the original equation didn't have logs I can still use them?

OpenStudy (anonymous):

@amistre64 thank you so much for providing the steps! I just want to make sure I got them correctly. Now I have a+8.5log14=log6... do i divide the logs?

OpenStudy (amistre64):

yes, the logs just represent some numbers, log14 has a constant value, and so does log6 ... so they are just numbers at this point

OpenStudy (anonymous):

so i would just do 14/6 which is 2.33333 and then subtract 8.5 from that to figure out a?

OpenStudy (amistre64):

logs are inverses of exponents ... just like subtraction is the inverse of addition. when you have exponents, you can use logs to undo things

OpenStudy (amistre64):

lets say log14 = s, and log6 = t (a+8.5)log14 = log6 (a+8.5)s = t a+8.5 = t/s t/s is not 6/14, it is log 6/ log 14, or simply:\[log_{14}(6)\]

OpenStudy (amistre64):

or another way to see it .. if this is to hard. is to use the property:\[n^{log_n(k)}=k\] \[14^{log_{14}(6)}=6\]therefore\[a+8.5=log_{14}(6)\]

OpenStudy (anonymous):

i see that makes more sense now that I see all of the little steps! so to figure out a from there i would solve logv14(6) and subtract 8.5 from whatever i get?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

would my answer be correct if i got a= -7.821?

OpenStudy (amistre64):

be careful that the solution is asking for an approximation, or an exact exact solution would keep it as \(a=log_{14}(6)-8.5\) approx solution would round \(log_{14}(6)-8.5\) to some number of decimal places

OpenStudy (anonymous):

it says to solve and then round my answer to the nearest ten-thousandth... so how would ido that?

OpenStudy (amistre64):

then approximation is fine http://www.wolframalpha.com/input/?i=log_%7B14%7D%286%29-8.5

OpenStudy (amistre64):

the nearest ten thousandths is 4 decies 10 000 ^^^^ 4 zeros, so 4th decimal place

OpenStudy (anonymous):

awesome! thank you so much for all of your help!! :)

OpenStudy (amistre64):

youre welcome